40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 52 of 55  Not logged in ELOG logo
ID Date Author Type Category Subject
  191   Wed Jul 22 20:36:37 2009 Aidan, ConnorLaserFiberAligned beam into fiber ... mode-matching needs a little work

Connor and I got the beam aligned into the fiber today. After a little optimization of the fiber coupled XYZ I'd estimate we were getting maybe 10-15% transmission. I'm pretty sure we can improve this with some work. We caculated the input mode (which I will reproduce elsewhere) but it's looking for a mode size of approximately 6.6 microns diameter. With a 10X microscope objective (f = 16.5mm) we want the intput beam to the coupler to be around 3.4mm diameter - which we have.

We mounted the AOM and inserted it into the setup. We should be able to get an 80MHz frequency shifted beam through the fiber tomorrow measure the fiber noise. That same setup can be used in a PLL to suppress the fiber noise.

 

  190   Wed Jul 22 11:09:45 2009 MichelleMisc Weekly Update 5

Over the course of the past week I've done a few things. When I began alignment I discovered that our laser was periodically shutting itself down. This was a very perplexing problem for about 2 days before Alberto came in with the diagnosis: the diode was overheating. We'll be sending that in to get it fixed, and we'll also put a heat sink on its casing. Hopefully that will be up and running by the time we really need it.

Right now, Aidan's laser is set up with a 50/50 beamsplitter, and then a half wave plate at each output of that. That way we can run the laser at full power, and each of us can independently adjust the power going into our respective experiments. This is working well while we're aligning things, but it clearly won't work long-term - we need to act directly on the laser's frequency to lock it to a cavity. This may not fare so well for fiber noise suppression.

We have our setup mostly aligned. The beam is going through some steering mirrors, through a lens, into the Faraday Isolator, through another set of steering mirros and a lens, and into the EOM. There is very little loss inside the isolator (putting in ~35 mW and getting out 33-34 mW), however I have not been able to get the power at the output of the EOM higher than ~ 24 mW. I don't think this is normal, but I will check that with people who know better than I do. I think it is probably the fault of poor alignment - the aperture is ~ 2mm and it's about 2 cm from the lens in the middle of the table, so it's hard to reach it or even view it properly to see what needs tweaking.

Over the next week I plan to finish the alignment and hopefully get a lock. I'll get a picture up once the rest of the setup is in place and aligned.

 

Oh, and I also helped clean the lab this week. It's pretty shiny, except for the heaps of garbage boxes now sitting in the hallway. We'll take care of that soon.

  189   Wed Jul 22 09:43:53 2009 ConnorMisc Weekly Report #5

Most of this week has been spent aligning optics for the fiber stabilization experiment.  Part of the setup is shown here

The next step is to couple the beam into the fiber, single pass it and interferometrically beat it with the original signal so we can characterize the acquired phase noise.

Mode matching to various optical elements involves changing the beam parameter (i.e., beam width and radius of curvature) to fit specifications. We do this using lenses. Given the incoming beam parameter, the desired outgoing beam parameter, and lens focal lengths, we can find solutions (lens positions) to the problem.

  188   Tue Jul 21 20:41:56 2009 DmassLaserPSLRazor Blades!

Quote:
Yeah, real micrometer with a reading is > plain screw.


Converting from flemish ells was hard. I'll just use a mic next time.
  187   Tue Jul 21 20:39:18 2009 DmassLaserPSLPSL Beam Scans

I took a bunch of scans of the PSL with the Beamscan and compared them with some razorblade occlusion measurements

The first steering mirror after the beam exits the PSL was removed, and data was taken at this point in the beam path.

The numbers for the waist derived from my razorblade measurements are on the plot in green. The disagreement seems not horrible to me.

 

We should still fix the beamscan, and think about buying another (newer) one.

 

I am told that the beamscan axis drawn on the head (labeled "1" and "2"), not the slit axis is what sets the direction of the field you are sampling. Data for both vertical and horizontal spatial modes of the beam are included. There is some ellipticity to the beam that I will have to account for in choosing my (new?) mode matching solution.

 

Error bars are included on the green point, they are just small.

Attachment 1: VertPSLScan.pdf
VertPSLScan.pdf
Attachment 2: HorPSLScan.pdf
HorPSLScan.pdf
  186   Tue Jul 21 18:45:26 2009 ConnorLaserFiberalignment of optics

The 1W NPRO is, for the moment, being used for both the FS and Gyro experiments. Aidan and I began aligning the optics for the FS experiment today.

The beam is first passed through a quarter-waveplate to linearize its polarization. A 50-50 beamsplitter picks off the beam for the Gyro experiment. The transmitted beam (for FS) is then attenuated by a half-waveplate and PBS. We steer the transmitted beam through a lens, a Faraday Isolator, and another lens to collimate it.

A photograph of the setup is attached.

We plan to direct the beam through an AOM and couple it to the fibre so we can characterize the acquired phase noise.

Attachment 1: FSprelimsetup.jpg
FSprelimsetup.jpg
  185   Tue Jul 21 02:09:45 2009 DmassLaserPSLPSL Realigned, now better

Quote:

Koji came down to the dungeon to help me with the "what is the junk coming out of the PSL" problem.

We noticed that the beam right after the PBS looked funky, and the beam before the power amplifier stage inside the PSL looked fine.

We decided to realign the optics inside the PSL. We only touched the two steering mirrors immediately preceeding the power amplifier.

We got a 35% increase in power and the output mode looked cleaner to the eye. (No more solar flares)

We are now happy.

 

 The following image contains approximations of the changes we made to the PSL steering mirrors (~5 degree accuracy)

Attachment 1: PSLRealignment.png
PSLRealignment.png
  184   Mon Jul 20 23:19:49 2009 ranaLaserPSLRazor Blades!
Yeah, real micrometer with a reading is > plain screw.
  183   Sun Jul 19 09:44:09 2009 AidanComputingGeneralAliases added for standard LIGO tools

 

I've updated the /home/controls/.bashrc file to include some aliases to standard ligo tools. These should now be accessible anywhere from the command line.

alias dtt="/bin/diaggui"
alias dmt="/bin/dmtviewer"
alias foton="/bin/foton"
alias dv="dataviewer"
 
StripTool is already accessible from the command line.

  182   Fri Jul 17 15:39:31 2009 DmassLaserPSLRazor Blades!

Quote:

Since the beamscan is in a questionable state of scanny goodness, Koji advised that I do some razor blade occlusion power measurements of the beam and then fit an erf to it to find the waist. I took data with the tinker toys pictured below.

I will compare these results with some beamscans results to verify (hopefully) that the beamscan is outputting useful results, not lies.

 Using the setups in the quoted post, I took manual beamscans 23" in front of the PSL enclosure, before the first steering mirror, and fit P = A erf(B*x + C)

 I measured the relative position of my razorblade with a micrometer and calculated the error from an estimated uncertainty of it's angle. This seemed to agree with repeatability of measurements for a given experimental state. Uncertainty < .002"

I watched the Power meter for ~ 60 seconds for each measurement, it fluctuated around some point and seemed to not be drifting @ DC, so the upper and lower error bars of each point included are the bounds of the fluctuation of the power meter. These were less than +/- 5mW about my points, so a fractional uncertainty of about 2% at my maximum power.

 

I got waists of:

Vertical: 799 +/- 4.5 microns

Horizontol 827 +/- 1.2 microns

Attached plot includes data w/ error and functional form of fit

 

As expected, my Chi^2 is "bad" since I am fitting the input beam to the PMC with a 00 mode description of the waist, which ignores all higher order modal content.

Attachment 1: RazorScans.pdf
RazorScans.pdf
  181   Thu Jul 16 23:21:40 2009 DmassLaserPSLPSL Diode Powers and temperatures

At Anamaria's request, I recorded the diode powers (and temperatures) of the PSL.

~5 min after power up they were:

temp/power

20C  / 29.8W

23.8C / 28.6W

24C / 25.2W

25C / 27.8W

6 hours later they were

20C / 29.0W

23.8C / 28.1W

24C / 24.8W

25C 27.3W

  180   Thu Jul 16 17:46:41 2009 DmassLaserPSLPSL Realigned, now better

Quote:

Koji came down to the dungeon to help me with the "what is the junk coming out of the PSL" problem.

We noticed that the beam right after the PBS looked funky, and the beam before the power amplifier stage inside the PSL looked fine.

We decided to realign the optics inside the PSL. We only touched the two steering mirrors immediately preceeding the power amplifier.

We got a 35% increase in power and the output mode looked cleaner to the eye. (No more solar flares)

We are now happy.

 

 

I will add pictures of what knobs we touched and how much each was turned by here

  179   Thu Jul 16 17:43:59 2009 DmassLaserPSLRazor Blades!

Quote:

Quote:

Since the beamscan is in a questionable state of scanny goodness, Koji advised that I do some razor blade occlusion power measurements of the beam and then fit an erf to it to find the waist. I took data with the tinker toys pictured below.

I will compare these results with some beamscans results to verify (hopefully) that the beamscan is outputting useful results, not lies.

 

Actually, it would be nicer if we have a calibrated micrometer screw for the thread.
Connor and I tried to make another set of the razor blade arrangement with a micrometer.
We put it on the PSL optical table. Please use it.

 

I got about a .002" error using a micrometer with my other setup, I expect this will suffice. If the errors in position end up too large I will use the new setup.

  178   Thu Jul 16 15:01:48 2009 Koji & Connor LaserPSLRazor Blades!

Quote:

Since the beamscan is in a questionable state of scanny goodness, Koji advised that I do some razor blade occlusion power measurements of the beam and then fit an erf to it to find the waist. I took data with the tinker toys pictured below.

I will compare these results with some beamscans results to verify (hopefully) that the beamscan is outputting useful results, not lies.

 

Actually, it would be nicer if we have a calibrated micrometer screw for the thread.
Connor and I tried to make another set of the razor blade arrangement with a micrometer.
We put it on the PSL optical table. Please use it.

  177   Thu Jul 16 04:59:38 2009 DmassLaserPSLRazor Blades!

Since the beamscan is in a questionable state of scanny goodness, Koji advised that I do some razor blade occlusion power measurements of the beam and then fit an erf to it to find the waist. I took data with the tinker toys pictured below.

I will compare these results with some beamscans results to verify (hopefully) that the beamscan is outputting useful results, not lies.

Attachment 1: Photo_1.jpg
Photo_1.jpg
Attachment 2: Photo_3.jpg
Photo_3.jpg
  176   Wed Jul 15 22:13:09 2009 ranaLaserGeneralGyro NPRO
I just came down and got the same behavior. With the diode current cranked up all the way, the output power was 700 mW.
This is good, since this is supposed to be a 700 mW NPRO.

I recommend that we pick a diode current (whatever we get with ADJ = 0) and stick with it. Use a PBS for changing the power and not the diode current.

In order to actually run the NPRO for extended periods of time we need a heatsink mounted on its side. Otherwise these things get too hot and the TEC for the diode can't keep up. I'll ping Peter King for one, but if he doesn't have one we'll just buy one from a catalog.

The lab temperature also seemed warm, like Michelle noted. The AC controller had the right setting so I think its that the valve in the overhead ducting has gone bad again. It failed twice in the first couple weeks that the lab was opened. I have replaced the battery in the digital temperature sensor and put it on the gyro table. After I lowered the temp. setpoint on the wall by a couple degrees it came down to 72 F. I also re-activated the circular paper temp. recorder and put it on the First Aid kit. Its in 24-hour mode so we can get a trend.

I will call up PMA and tell them we need a temperature consult again.


Also, the lab is not your personal trash receptacle. Please don't store empty cardboard boxes or drinks in there. Put tools back where they go when you're done and don't leave a pile of booties and hats laying on the floor.
  175   Wed Jul 15 16:05:53 2009 DmassLaserPSLPSL Realigned, now better

Koji came down to the dungeon to help me with the "what is the junk coming out of the PSL" problem.

We noticed that the beam right after the PBS looked funky, and the beam before the power amplifier stage inside the PSL looked fine.

We decided to realign the optics inside the PSL. We only touched the two steering mirrors immediately preceeding the power amplifier.

We got a 35% increase in power and the output mode looked cleaner to the eye. (No more solar flares)

We are now happy.

 

  174   Wed Jul 15 14:40:44 2009 MichelleLaser NPRO ... fixed?

Alberto came over to have a look at the laser and discovered that the diode temperature was continuously increasing the longer the laser was kept on. The crystal temperature remained constant. He turned off the laser and shut the shutter, jiggled the cable connected to the laser a bit, and turned it back on. Lo! The temperature no longer increased...

So for the time being the laser seems to be fixed. If these problems start happening again, however, I may have to do some more rigorous troubleshooting/actually find out what's causing the problem.

  173   Wed Jul 15 12:36:17 2009 Connor MooneyMiscGeneralWeekly Report #4

     Aidan was in Hanford this week, so I switched gears and worked mostly on modematching the 35W PSL with Dmass.  This requires passing the beam through a couple of lenses to alter the waist size and location before it enters the PMC. Last year, he took some beamscans post-lenses. We wanted to see if this data matched with the pre-lens beam parameter. I used ray transfer matrices to perform the inverse calculation. We found a mismatch between the calculated and expected waist location, which could not be fully explained by the engineering tolerance of the lenses or by the lens orientation.

     Dmass also noticed something funny about the head of one of the lenses, so he made some changes to that and made more beamscans that we will look at.

     There are a number of objectives for the upcoming week:

     1) We believe that the 1W NPRO is elliptically polarized. I learned about Jones vectors and Jones matrices, which are a convenient linear-algebraic representation of a beam's polarization and how it changes as it passes through various optics. The beam can be linearly polarized by passing it through a quarter-waveplate (QWP) whose fast axis is correctly angled. To test this, I will pass the beam through a QWP, and aligned a half-waveplate (HWP) and polarizing beamsplitter (PBS) to find its extinction ratio for various QWP angles. (The extinction ratio is the ratio of the minimum to maximum transmitted powers of the beam as the HWP is rotated.)

     2) I'll work on coupling the 1W NPRO into a 50m fiber.

     3)I'll align a Mach-Zehnder interferometer to measure the phase noise of the 1W NPRO, and also single pass the beam in one arm through the 50m fiber and characterize the phase noise it acquires.

    

 

  172   Wed Jul 15 10:47:55 2009 MichelleLaserGeneralNPRO acting up

Quote:

I'm having some issues with the 495 mW NPRO, or possibly with its power supply. I ran the laser for about 15 minutes at 2 mW while doing some aligning. I closed the shutter to move a mirror, and upon opening the shutter the laser was no longer lasing. The power supply read 0.0 A driving current. I closed the shutter again and checked all connections between the laser and the power supply. All in order. I opened the shutter and turned the dial a bit on the power supply, and lo and behold the current was back up to 0.46 A, the lasing threshold, but still no light coming out.

The drive current jumped around a lot, sometimes to 0.0 A, and even when I put it up to around 0.54 A, which the laser should definitely be lasing at, nothing came out the aperture. The internal laser temperature was normal and not fluctuating. The lab itself is pretty warm today, in the 70s. Unless someone changed the temperature, it felt like the heat was on. I don't think this would have any effect on the laser, but I'm not really sure.

So, I'm going to search for the manual online and see if anyone over at the 40m had problems with this power supply.

 I took the power supply over to the 40m to see if Alberto's laser would work when hooked up to it. It did, so I suspect that the problem lies with the laser rather than the power supply. Alberto was able to get double the drive current that we were getting, so that is also suspect. He said the "ADJ" readout (power adjustment) on the supply was usually set to zero, but when the "Pwr" readout (supposedly the power coming out of the laser) is changed, it changes the ADJ and drive current and anything else which would obviously be associated with power. I hooked up the power supply to my laser again and here's what I found:

-It lases for short periods of time, usually <5 mins, before it stops.

-It lases for longer periods of time at higher drive currents.

-I still can't get more than 1.02 A and less than 0.30 A out of the power supply.

-The "Pwr" display on the supply still doesn't correspond to the power as measured by the Ophir powermeter.

My next steps are to see if Alberto can come and take a look, in case I'm doing something on the operator side which is obviously wrong. I will also consult with Rana.

  171   Tue Jul 14 15:08:07 2009 Connor MooneyLaserPSLBeam Divergence of PSL

A while ago, Dmass made some beamscans of the PSL beam after it passed through two lenses. Here is a link to the Dmass Modematching elog.

We wanted to make sure that these measurements are consistent with what we think the incoming beam parameter is.  I wrote a matlab file which works backwards from his data to find the incoming beam parameter using ABCD matrices.  I used the focal lengths indicated on the lenses.

I found a disagreement of 20cm in the PSL waist location with a yesteryear elog entry  Dmass checked to see if my computation matched with his, and got the same result.

The first thing that we thought might have been wrong is that the focal lengths need to be switched, but switching them didn't significantly our change our results (in fact, it made them worse).

We also suspected that the engineering tolerance of the lenses might allow for the difference. I tried changing the focal lengths of the lenses by 10% for both lens orientations, but the computations made only small changes to our original result, so this is probably not the culprit.

 

 

 

  170   Mon Jul 13 21:39:11 2009 MichelleLaserGeneralNPRO acting up

Quote:

I'm having some issues with the 495 mW NPRO, or possibly with its power supply. I ran the laser for about 15 minutes at 2 mW while doing some aligning. I closed the shutter to move a mirror, and upon opening the shutter the laser was no longer lasing. The power supply read 0.0 A driving current. I closed the shutter again and checked all connections between the laser and the power supply. All in order. I opened the shutter and turned the dial a bit on the power supply, and lo and behold the current was back up to 0.46 A, the lasing threshold, but still no light coming out.

The drive current jumped around a lot, sometimes to 0.0 A, and even when I put it up to around 0.54 A, which the laser should definitely be lasing at, nothing came out the aperture. The internal laser temperature was normal and not fluctuating. The lab itself is pretty warm today, in the 70s. Unless someone changed the temperature, it felt like the heat was on. I don't think this would have any effect on the laser, but I'm not really sure.

So, I'm going to search for the manual online and see if anyone over at the 40m had problems with this power supply.

 We should get back to trending the lab temperature. I think there are some long standing problems with the HVAC setup. It would help to quantify this.

  169   Mon Jul 13 13:59:19 2009 MichelleLaserGeneralNPRO acting up

I'm having some issues with the 495 mW NPRO, or possibly with its power supply. I ran the laser for about 15 minutes at 2 mW while doing some aligning. I closed the shutter to move a mirror, and upon opening the shutter the laser was no longer lasing. The power supply read 0.0 A driving current. I closed the shutter again and checked all connections between the laser and the power supply. All in order. I opened the shutter and turned the dial a bit on the power supply, and lo and behold the current was back up to 0.46 A, the lasing threshold, but still no light coming out.

The drive current jumped around a lot, sometimes to 0.0 A, and even when I put it up to around 0.54 A, which the laser should definitely be lasing at, nothing came out the aperture. The internal laser temperature was normal and not fluctuating. The lab itself is pretty warm today, in the 70s. Unless someone changed the temperature, it felt like the heat was on. I don't think this would have any effect on the laser, but I'm not really sure.

So, I'm going to search for the manual online and see if anyone over at the 40m had problems with this power supply.

  168   Thu Jul 9 13:29:08 2009 AidanComputingDAQMore from Alex on frame builder disk being full

 Hey Alex,


I've got a couple of questions (just for my own understanding).

- I thought the frames were supposed to overwrite old data. Why then  
did the disk appear full? And is it likely to appear full in future?  
If so how do we deal with it?

I reduced the total number of data directories from 120 to 60. I think the
frame size has increased and  so the disk filled up. daqd program, the
version you have there, erases old frames based on total number of
directories, it doesn't look at the disk full percentage. I have a new
version which doesn't do that, so we may want to install that new version
at some point, especially when you guys decide to upgrade to the latest
real-time code generator software. The new version of daqd/nds is not
erasing anything but simply relies on a cron job to erase old files.

- why did you reduce the number of directories in /frame/full to 60?

Thanks for your help,
Aidan.
 

  167   Thu Jul 9 10:48:03 2009 ConnorLaserFiberBeamscan for 1w NPRO

Aidan suggested that we try to characterize the beam width of the 1W NPRO again, this time operating the laser at high power and attenuating the beam with a half-wave plate and a polarizing beam splitter.  I aligned the optics such that at full output power (~1.19W), the power incident on the BeamScan is 145 mW.

There are a number of things I noticed about the beam profile around 10" from the laser aperture that I don't know how to deal with, so I decided not to make measurements until working things out.

While the profile of the beam is much cleaner now than it was before when we measured beam width at low power output, the width of the beam seems to depend on the driving current to the laser.  At 1.2A driving current, the beam diameter was roughly 1650 micrometers.  The width dropped down to around 1300 micrometers at 2A driving current, and then climbed back up to 1500 micrometers at full power output with 2.38A driving current.

Also, as David noted, the beam width fluctuated by up to 40 micrometers when moved to different places on the head.

  166   Thu Jul 9 10:46:52 2009 AidanComputingDAQDeleted old /frame/full directories and rebuilt directory structure ...

 

This was also from Alex:

 

Too late, I deleted /frame/full data already, sorry.

So I changed the number of full frames dirs from 120 down to 60 and
restarted the frame builder and c2atf front-end. Looks like the data is
written out to disk now. Please check it out.

Alex

  165   Thu Jul 9 10:45:16 2009 AidanComputingDAQAlex's frame builder problem diagnosis ...

 

This from Alex: 

 

The filesystem is full. If you look at the log messages:

[controls@oms fb]$ pwd
/cvs/cds/caltech/target/fb
[controls@oms fb]$ tail -5 logs/daqd.log.19816 
[Thu Jul  9 10:07:51 2009] Couldn't open full frame file
`/frames/full/data38/.C-R-931194464-16.gwf' for writing; errno 28
[Thu Jul  9 10:07:51 2009] Couldn't open full trend frame file
`/frames/trend/second/data2/C-T-931194420-60.gwf' for writing; errno 28
[Thu Jul  9 10:08:07 2009] Couldn't open full frame file
`/frames/full/data38/.C-R-931194480-16.gwf' for writing; errno 28
[Thu Jul  9 10:08:23 2009] Couldn't open full frame file
`/frames/full/data38/.C-R-931194496-16.gwf' for writing; errno 28
[Thu Jul  9 10:08:39 2009] Couldn't open full frame file
`/frames/full/data38/.C-R-931194512-16.gwf' for writing; errno 28
[controls@oms fb]$ df -h /frames/
Filesystem            Size  Used Avail Use% Mounted on
/dev/mapper/VolGroup00-LogVol00
                     275G  261G     0 100% /frames


I stopped the frame builder for now until we figure out the data
directories. Can I delete all the data in /frames/full and recreate a new
set of directories? Otherwise this is too messy to rearrange. Let me know.

-alex


On Wed, 8 Jul 2009, Aidan Brooks wrote:


Hi Alex,

I fixed the problem with being able to telnet into fb0. There was just a typo in the file:
/cvs/cds/caltech/chans/daq/C2ATF.ini
Once I fixed that the frame builder rebooted just fine.

The problem at the moment is that the data does not appear to be writing to file. If I rebuild the front end from the start, so I go to oms and run the following commands:

cd /cvs/cds/advLigo
make realclean
make atf
make install-atf
make install-daq-atf
make install-screens-atf

And if I edit the file C2ATF.ini so that a single channel, e.g. C2:ATF-CTRL_PMC_REFL_IN1_DAQ is set to acquire. 

and killatf, followed by startatf and reboot the frame builder

Then: the front end starts okay and I can see signals in the EPICS screen that make sense. I can run StripTool and see the real time time series of a channel, I can start dataviewer and see C2:ATF-CTRL_PMC_REFL_IN1_DAQ in the list of fast channels. I can run the realtime viewer in dataviewer and see that channel realtime in a Grace window. However, I cannot get a playback of that channel to work and if I run dtt I can only get a spectrum if I specify a start time for that spectrum within the last (approximately) 30s in the past.

The data rate in the .mdl simulink model is 16K, the data rate in the C2ATF.ini files is 16K, the data rate in the file:
/cvs/cds/caltech/target/fb/daqdrc is set to 16K and there are 120 data directories in the /frames/full directory.

I can't think of any reason why the data is not being written to file. Can you log in and figure out what's wrong? You should be able to get access via controls@131.215.114.183

Thanks for your help,
Aidan.
 

  164   Thu Jul 9 02:24:07 2009 ranaLaserPSLMode Matching and the PMC
The Anderson formulas are sort of OK in this regime but should only be used to estimate the mismatch.
Of course, its easy just to do the overlap integral and get the exact number. Its of limited use, since
the cavity will often not lock at the right point of the mismatch is so big.

Another possible method is to turn the power way down (less than 1 mW) and then scan it and fit
for the modes in transmission. The low power is to avoid thermal self-locking distortion of the fringe.

One can use knowledge of the PZT coefficient and the wavelength of the laser and modulation frequencies
to determine the cavity's g-factors and therefore the waist sizes. The cavity g-factor is defined as
g = 1 - L/R.
  163   Wed Jul 8 19:05:49 2009 DmassLaserPSLMode Matching and the PMC

I measured the waist of the input beam to the PMC after putting the mode matching lenses back. In the process of this, I noticed some problems with the beamscan.

I played with the beamscan axis until I found something that looked gaussian to the eye, and took a series of measurements, and fit the data (see attached).

 

 

I calculated the expected power loss from mode mismatch a la Dana Anderson's paper on cavity (mis)alignment and got some results:

I get estimates for power losses of:

6% due to waist position mismatch

12% due to waist size mismatch

NOTA BENE: The formulas in the Anderson paper start to break down as the linear approximations break down, which is happening for us...

 

If b is my mismatch parameter in waist position, my condition for the linear approximation is:

lambda*b/(pi*waist^2) <<1          but the LHS quantity is .492

and if w0' and w0 are my input beam and cavity waists, respectively, my condition is:

w0'/w0 - 1 <<1                            but the LHS is .34 here

Summary: I don't know how bad the Anderson formulas are in this regime of mismatch, but I can clearly do better in my mode matching. Next step: mode match/fix lens positions

 

Attached Image:

RED is roughly the PMC waist and position (based on the 40m PMC's waist - we have no measurement of this PMC's waist size yet afaik: put this on to do list)

GREEN is weighted least squares fit

BLUE is the data

 

Attachment 1: PMCbeamwaist.pdf
PMCbeamwaist.pdf
  162   Wed Jul 8 17:29:46 2009 Connor Mooney, Michelle StephensLaserFiberBeam Divergence of 495mW NPRO

We made a set of beam width measurements of the 495 mW laser using the Beamscan. All values were taken at (1/e^2) of the peak intensity, in micrometers. The distances were taken in units of inches (no metric-based rulers).

Here is our data: 

Distances (in inches): [2 3 4 5 6 7 8 9 10 12 14 16 18 20 22]

Beam Diameters (in micrometers): [357 364 413 448 528 608 690 758 854 1043 1216 1403 1624 1812 1999]

Again, we least-squared fit the data to the characteristic curve parametrized by the waist location and size, and found that the waist size is 178 micrometers, located 4.6 cm. in front of the aperture.

THINGS TO NOTE:

First, DMass mentioned he suspects that the beamscan is impaired, so we're not sure how well it is acting.

Second, Rana noted that the best indicators of the parameters of interest are data points taken far from the waist, where the beam width grows close to linearly with distance.  We took  lots of data points in the nonlinear regime, as seen on the attached plot. It might be a good idea to take more data points further away and refit our data.

Lastly, this laser is more stable than the 1W NPRO, with fluctuations of only 1-3 micrometers around the reported beam width values rather than the 40-50 we saw for the 1W.

The directory for our code is /users/cmooney/Beamscan_495mW.m

Here is the code:

%This is a graphical representation of beam width data taken for the 495 mW NPRO
%Laser, taken on July 8, 2009.

%W1 is the full width at 1/(e^2)*(peak intensity) taken at 1.00A driving
%current (around 150 mW) for the axis A1 as labelled on the BeamScan. Measurements are
%in micrometers.

W1 = [357 364 413 448 528 608 690 758 854 1043 1216 1403 1624 1812 1999];

%S1 is the standard deviation of the measurements given in W1.

S1 = [0.3 0.3 0.4 0.4 0.4 1.2 0.8 0.8 1 1 2.7 2.9 1.6 1.6 2.2];

d = 2; %distance from laser aperture to its base, from which it is easier to measure distance
D = [0 1 2 3 4 5 6 7 8 10 12 14 16 18 20]; %distances (in inches) at which measurements were taken
Ds = 2.54*(D+2)/100; %conversion to meters

W1s = 0.5*W1/(10^6);
S1s = S1/(10^6);

errorbar(Ds,W1s,S1s,'.')
title('Beam Width vs. Distance from Laser');
xlabel('Distance (m)');
ylabel('Beam Width (m)');

%This is a brief program to assist in the measurement of width of a
%Gaussian beam. Its wavelength is 1064 nm.

xdata = Ds;
ydata = W1s;
 
a0 = [.0001, .2]  % Starting guess
[a] = lsqcurvefit(@myfun,a0,xdata,ydata)

w_0 = a(1);
Zo = a(2)
L = 1064*10^(-9);
Dmax = 1;
sh = 0.35;
dt = (Dmax+sh)/10000;
Z = -sh:dt:Dmax;

Wt = w_0*(1+(((Z+Zo).*L)/(pi*w_0^2)).^2).^(1/2);

hold on;
plot(Z,Wt);

Attachment 1: beamdiv495mW.pdf
beamdiv495mW.pdf
  160   Wed Jul 8 10:46:58 2009 ConnorMiscGeneralWeekly Report #3

          The majority of this week has been spent learning about sources of intensity noise and the methods we use to characterize it.  The sources that I focused on were the "electronics noise" associated with various instruments (this week the instrument used was a photodetector, contributing fluctuations in intensity called "dark noise"), shot noise, and the noise produced by source itself (in our case, a 495 mW NPRO Laser).

          The theoretical method of characterizing a noisy time series is called a Power Spectrum, which is defined as the Fourier transform of the series' autocorrelation.  Intuitively, this is a frequency-domain representation of the extent to which a given frequency is represented in the time series.  Experimentally, it is more convenient to take advantage of a couple of theorems relating the Power Spectrum of a time series to its Fourier transform over a finite time interval;  The SRS spectrum analyzer that we used this week calculates the discrete Fourier transform of an input signal over a given time interval, squares that value, and divides by the time interval to get an estimation of the Power Spectrum.  Typically, we run the signal into the spectrum analyzer for many of these time intervals, and the values computed over each of these intervals are averaged.

          We measured the dark noise and the total intensity noise of a 495 mW NPRO laser using a PDA10CS photodetector, and made a theoretical prediction for shot noise based on the output power of the laser.  The data we collected is contained in yesterday's entry.

          Aidan has been working hard on getting the DAQ to its operating point, which will make measurements such as those described above easier to make.

          The plan of action for the upcoming days is to set up a Mach-Zehnder Interferometer for the same laser and characterize the phase noise its beam acquires, hopefully with the assistance of the DAQ.

 

  159   Wed Jul 8 10:36:44 2009 MichelleMisc Week 3 Update

We have mirrors! And clamps! And posts! In short, everything we need to actually put all the optics in their proper places. Thus, this week I have successfully placed the steering mirrors and the high reflectance mirrors into their mounts without touching/breaking any of them, and soon I'll be clamping them in their proper locations.

Connor and I  found out that Aidan's laser isn't behaving itself. The manufacturer specifies the waist as being in a different location than where we calculated it to be. We sent the beam through a half-wave plate and then a polarizing beam splitter and then measured the power along each axis, and the ratios are strange. We can't get the beam to be entirely transmitted along one axis by rotating the half-wave plate, like we should be able to if the beam is coming out linearly polarized. So the hypothesis is...it's not coming out linearly polarized. It may be elliptical. For the time being we're going to split the beam from the other NPRO and use it for both experiments, at least while the alignment is taking place. In the case of the gyro, we'll actually be feeding back to the laser and altering its frequency, so we'll definitely need separate lasers for each experiment at that time.

The 495 mW NPRO's SOP has finally been (mostly) approved and it first lased on Monday. This laser's power supply is a bit weird - the display for the power it thinks the laser is putting out doesn't match at all with what the powermeter actually says the laser is putting out. So we're not paying attention to the power display. The drive current goes up to 1.00 A, and at this maximum current, the laser is putting out ~150 mW. Apart from not being able to get a higher power than this, the laser is fairly well behaved - the power fluctuations as measured by the powermeter are small. The results of output power vs. drive currentmeasurements are on the eLog.

Connor and I also measured the electronics noise and intensity noise at various different powers, and then did a theoretical shot noise calculation for each of these. The setup and procedure are described in more detail in a recent eLog entry, which also includes graphs and the matlab code used to generate them.

Over the next week I'll be setting up the optics table for a simple PDH lock using a fabry-perot cavity, aligning the laser, and setting up the proper feedback using the DAQ. (Which I'll need to talk to Dmass about).

Oh, and we found out yesterday that Connor and I can't do any hazardous tasks in the lab without supervision, which includes aligning the lasers, even at low power. Aidan and Alastair will both be gone from tomorrow until next week Friday, so this poses a slight challenge. Hopefully Dmass will be around?

That's about it for now, I think.

  158   Tue Jul 7 16:49:05 2009 Michelle Stephens, Connor MooneyLaserGeneralNoise Measurements

We used an SRS Spectrum Analyzer in parallel with an oscilloscope to measure various sources of noise in the 495 mW NPRO laser. We sent the beam into a PDA10CS photodetector and fed that signal to the spec analyzer and the scope. We measured the dark noise, and then the total noise at 2 mW, 4 mW, 7mW, and 9mW. We then did a theoretical calculation for the shot noise, as follows:

N = (Po*dt) / (h*f)

dN = sqrt(Po*dt / h*f)

dE = sqrt(h*f*Po*dt)

dI = dE*(Responsivity of PD)

dV = dI*(Gain of PD)

We plotted all three of these noise sources in matlab. The plot and the code are attached. The directories are:

/users/mstephens/NPRONoise.m

/users/mstephens/495mW_NPRO_Noise.pdf

 

Attachment 1: 495mW_NPRO_Noise.pdf
495mW_NPRO_Noise.pdf
Attachment 2: NPRONoise.m
load -mat SRS003.MAT
load -mat SRS005.MAT
load -mat SRS006.MAT
load -mat SRS007.MAT
load -mat SRS008.MAT

h = 6.626E-34;
mu = 3E8/1.604E-6;
deltaT = 1.0;

... 94 more lines ...
  157   Tue Jul 7 15:16:35 2009 AidanComputingDAQFrame builder not restarting ... fixed!

Fixed!

The problem was a typo that had cropped up in /cvs/cds/caltech/chans/daq/C2ATF.ini when I was editing this file. I had to fix that file and reboot oms to get the frame builder working again. Now it appears to be fine - I can telnet into fb0 again.

 

Quote:

 

I tried to reboot the frame builder last night, as below, and now it won't restart and I can't log into it. If I can find the box then I can try to physically reset it.

[controls@ws1 ~]$ telnet fb0 8087
dadq> shutdown

 

 

 

  156   Tue Jul 7 09:26:48 2009 AidanComputingDAQFrame builder not restarting ...

 

I tried to reboot the frame builder last night, as below, and now it won't restart and I can't log into it. If I can find the box then I can try to physically reset it.

[controls@ws1 ~]$ telnet fb0 8087
dadq> shutdown
  155   Mon Jul 6 21:03:47 2009 DmassLaserFiberNPRO beam width divergence measurements

Quote:

We used the BeamScan to characterize the divergence of the NPRO laser as a function of distance. Measurements were taken at (1/e^2)*Imax. We then fit the data to the function characteristic of gaussian divergence to find the location and size of the waist. The results are shown graphically in the attached file. Extrapolating, we found that the waist is 21.5 cm behind the laser aperture, and the (virtual, assuming that the laser beam encounters a diverging lens before the aperture) waist size is 124 micrometers.

 

Here is our data:

The distances, measured in inches from the base of the laser, are

[0 0.5 1 3 5 7 9 11 13 15 17 19 21 23 25]. The distance from the aperture of the laser to the base is 0.75 inches.

The beam sizes, measured in micrometers along one axis, are

[1334 1399 1542 1655 1968 2234 2537 2785 3063 3359 3638 4001 4253 4561 4708]

 

The directories for each of the Matlab files used to generate a graph of the data and carry out a least-squares regression are

/users/cmooney/BeamScans.m

/users/cmooney/myfun.m

 In general, zip and upload your files and matlab code. You or someone like you may want it later, and may really want to be able to get it remotely.

  154   Mon Jul 6 16:59:47 2009 Michelle Stephens, Connor MooneyLaser Power Output of 495 mW NPRO

We characterized the power output vs. drive current of the 495 mW NPRO laser for the gyro experiment. The current supply  goes up to 1.00 A. Here is our data:

C = [0.46 0.52 0.60 0.66 0.74 0.80 0.88 0.94 1.00];
Pd = [2 18 38 55 71 93 110 125 146];
Pu = [2 20 39 56 72 94 110 126 147];

C is drive current, Pd is lower limit on power output in mW, and Pu is upper limit.

We graphed the results and fit a line to it. The slope is 261.5 mW/A, and the intercept is -118.2 mW. The graph is attached.

directory is: \users\cmooney\Power495mW.m

Attachment 1: PowervCurrent495mW.pdf
PowervCurrent495mW.pdf
  153   Mon Jul 6 15:31:07 2009 Aidan, AlastairMiscGeneralEquipment moved from 58C

The following equipment was from 58C (old mesa beam lab) into 58D during the clean out of the former lab.

2x SR560s ([Ligo Project Property (LPP) C20862, Caltech 44961], [LPP: n/a, Caltech 44782])

2x Phillips PM 5193 programmable function generators (0.1 - 50MHz): ([LPP: n/a, Caltech 9313], [LPP: C20365, Caltech: 1348])

1x SRS DS345 function generator (<30MHz): (no tag - serial # 36560)

1x Tektronix AFG3102: (no tag - serial# AFG3102 C020795)

1x HP 8656B signal generator 1-990MHz: (LPP: n/a, Caltech 23616)

These have all been labelled "John Miller Lab" to indicate where they came from. (None of the equipment was plugged in)

  152   Sat Jul 4 12:42:06 2009 Aidan ComputingDAQDAq Map screen complete - just need to work out how to generate it automatically now

I've finalized the DAQ_MAP screen. All the current channels are accounted for. I've added some standard LIGO medm screen paraphenalia, such as the information button and the title bar.

I'd like to incorporate this into the build of a front end so that this screen is generated automatically. I'm a bit stumped as to how to get the LOST/SAVED indicators to work automatically. Maybe we can write something that will run in the startup of fb0.

A couple of extra notes:

1. The names of the channels in the frame builder do not exactly match the names in EPICS. This makes it impossible to drag and drop channel names into data viewer. Example:

C2:ATF-QPD1_SEG1 is the  channel defined in Simulink. C2:ATF-QPD1_SEG1_INMON is the channel as it appears in EPICS and C2:ATF-QPD1_SEG1_IN1_32768 is the channel as it appears in the frame builder. Can we add a channel in the frame builder that has the same name as the corresponding EPICS channel?

2. There seems to be a problem with CH32. It's working, in the sense that the channel exists, but when I plug the signal generator into the corresponding BNC connector on the DAQ I don't see any response on EPICS. Need to triple check that I haven't done made a typo anywhere.

Attachment 1: C2ATF_DAQ_MAP_v1.11.png
C2ATF_DAQ_MAP_v1.11.png
  151   Fri Jul 3 21:26:13 2009 AidanComputingDAQDAQ Map medm screen - added channel frame acquisition status

Quote:

I've created an medm screen that displays the current data channels associated with each of the connectors on the front of the DAQ. Basically its supposed to make it really easy to figure out what each channel in the DAQ physically corresponds to. I've added a link (DAQ MAP button) to it on the main medm screen

Eventually this should be automatically generated every time we rebuild the front end. I'd also like to include an indication on there of the units of each channel and whether they are in the frame or not.

medm screen: /cvs/cds/caltech/medm/c2/atf/C2ATF_DAQ_MAP.adl

Please do not edit this screen in the medm editor. I've added some comments to the file and moved the definition of each of the channels to the end of the file. The comments get erased when the screen is edited and the channel definitions get moved back to different places in the file.

 

 

I've added some indicators to the DAQ_MAP medm that show whether channels are saved to frame ("SAVED") or not ("LOST"). I'll have to add a routine to the frame builder start up so that these are updated whenever the frame builder is rebooted.

Attachment 1: C2ATF_DAQ_MAP_v1.1.png
C2ATF_DAQ_MAP_v1.1.png
  150   Fri Jul 3 19:49:30 2009 AidanComputingDAQDAQ Map medm screen - Connectors and Channel names

I've created an medm screen that displays the current data channels associated with each of the connectors on the front of the DAQ. Basically its supposed to make it really easy to figure out what each channel in the DAQ physically corresponds to. I've added a link (DAQ MAP button) to it on the main medm screen

Eventually this should be automatically generated every time we rebuild the front end. I'd also like to include an indication on there of the units of each channel and whether they are in the frame or not.

medm screen: /cvs/cds/caltech/medm/c2/atf/C2ATF_DAQ_MAP.adl

Please do not edit this screen in the medm editor. I've added some comments to the file and moved the definition of each of the channels to the end of the file. The comments get erased when the screen is edited and the channel definitions get moved back to different places in the file.

 

Attachment 1: C2ATF_DAQ_MAP.png
C2ATF_DAQ_MAP.png
  149   Thu Jul 2 21:40:57 2009 Aidan, ConnorComputingDAQFrame builder shenanigans

There appears to be something screwy going on with the frame builder - have some things been compiled and others not? Either that or I don't understand the system well enough.

We plugged a signal generator output a 0.1Hz sine wave into CH9 of the DAQ. This showed up in Alastair's medm Gyro screen (C2ATF_GENCONT.adl) in channel C2:ATF-GENERIC_GEN1_INMON. When we want this written to the frame, however, I have to uncomment the channel C2:ATF-CTRL_SIG_1_IN1_32678 in /cvs/cds/caltech/chans/daq/C2ATF.ini. This channel shows up in Dataviewer and the 0.1Hz signal is clearly visible.

I suspect that things have been only partially compiled recently and we need a clean rebuild of the front end and daq and frame builder. Anyway, we can get signals at the moment so we're actually going to do something useful, like characterize the intensity noise of the NPRO.

 

 

  148   Thu Jul 2 15:37:57 2009 AidanMiscFiberWeekly Report - query on previous reply

Quote:

Quote:
Most of this week was spent on three tasks: characterizing the beam divergence of the NPRO Laser and fitting a curve to the data, the progress report, and contending with serious computer problems. The information on the beam divergence can be accessed in a previous log. The waist size is around 120 micrometers, and according to a least-squares fit, it is located 21cm behind the aperture. My understanding of the fiber suppression setup and its relationship to the second-harmonic generation system was strengthened, but nowhere near completed, in the process of writing the progress report. We plan to send the NPRO beam through an AOM driven at f1, then double-pass it through a 50m fiber (not to be bent at a 90 degree angle as drawn in the meeting today) and another AOM driven at f2. This signal will be interferometrically mixed with the original beam from the NPRO, producing beats at 2(f2-f1), then run through a multiplying mixer driven at 2(f1-f2) and a low-pass filter so the final signal consists mostly of the noise acquired when travelling through the fiber. A computer characterizes this noise, and sends a signal to the first AOM which actuates on the original beam so as to suppress the noise. ********************NOTE: Aidan plans on changing the setup slightly, and I'll give the details on that in the next report************************* Finally, my computer acquired a terrible virus yesterday morning when I clicked on an innocuous-looking facebook message that was sent to my email and titled "hi." All of yesterday and the first part of today was spent rebuilding the computer. Don't trust facebook. I now have a centos operating system, so I'll be learning to use Linux fast (hopefully).

 FORMAT PLEASE!

 

Who was the author of the previous entry? It wasn't Connor apparently.

  147   Wed Jul 1 17:06:30 2009 MStephensMiscGeneralNew Gyro Design and Week 2 Summary

I've done a few things this week. Connor and I characterized his laser's beam divergence, the results of which can be found on a previous entry. I'll do the same measurement for my laser once we're allowed to turn it on. The table now has a mock-up of the Fabry-Perot cavity locking scheme, it's still missing a few mirrors but the gist of it is there. I also changed the setup for our actual gyro design. In order to increase the finesse of the cavity, we got rid of one of the partially transmitting mirrors in the triangular cavity. So there is now a setup with two highly reflective mirrors and one partially transmitting, and both the clockwise and the counterclockwise beams will be injected into that one. We're also going to set it up so the clockwise mode goes through the AOM twice in order to double the effect. My first progress report is attached to my 40m wiki page. (Sorry, no LaTeX yet, will figure that out).

Attached here is the new diagram for the gyro setup.

 

Attachment 1: NewerGyroDesign.PNG
NewerGyroDesign.PNG
  146   Wed Jul 1 16:37:33 2009 Connor MooneyMiscFiberWeekly Report

Quote:
Most of this week was spent on three tasks: characterizing the beam divergence of the NPRO Laser and fitting a curve to the data, the progress report, and contending with serious computer problems. The information on the beam divergence can be accessed in a previous log. The waist size is around 120 micrometers, and according to a least-squares fit, it is located 21cm behind the aperture. My understanding of the fiber suppression setup and its relationship to the second-harmonic generation system was strengthened, but nowhere near completed, in the process of writing the progress report. We plan to send the NPRO beam through an AOM driven at f1, then double-pass it through a 50m fiber (not to be bent at a 90 degree angle as drawn in the meeting today) and another AOM driven at f2. This signal will be interferometrically mixed with the original beam from the NPRO, producing beats at 2(f2-f1), then run through a multiplying mixer driven at 2(f1-f2) and a low-pass filter so the final signal consists mostly of the noise acquired when travelling through the fiber. A computer characterizes this noise, and sends a signal to the first AOM which actuates on the original beam so as to suppress the noise. ********************NOTE: Aidan plans on changing the setup slightly, and I'll give the details on that in the next report************************* Finally, my computer acquired a terrible virus yesterday morning when I clicked on an innocuous-looking facebook message that was sent to my email and titled "hi." All of yesterday and the first part of today was spent rebuilding the computer. Don't trust facebook. I now have a centos operating system, so I'll be learning to use Linux fast (hopefully).

 FORMAT PLEASE!

  145   Wed Jul 1 15:34:26 2009 Connor MooneyMiscFiberWeekly Report
Most of this week was spent on three tasks: characterizing the beam divergence of the NPRO Laser and fitting a curve to the data, the progress report, and contending with serious computer problems. The information on the beam divergence can be accessed in a previous log. The waist size is around 120 micrometers, and according to a least-squares fit, it is located 21cm behind the aperture. My understanding of the fiber suppression setup and its relationship to the second-harmonic generation system was strengthened, but nowhere near completed, in the process of writing the progress report. We plan to send the NPRO beam through an AOM driven at f1, then double-pass it through a 50m fiber (not to be bent at a 90 degree angle as drawn in the meeting today) and another AOM driven at f2. This signal will be interferometrically mixed with the original beam from the NPRO, producing beats at 2(f2-f1), then run through a multiplying mixer driven at 2(f1-f2) and a low-pass filter so the final signal consists mostly of the noise acquired when travelling through the fiber. A computer characterizes this noise, and sends a signal to the first AOM which actuates on the original beam so as to suppress the noise. ********************NOTE: Aidan plans on changing the setup slightly, and I'll give the details on that in the next report************************* Finally, my computer acquired a terrible virus yesterday morning when I clicked on an innocuous-looking facebook message that was sent to my email and titled "hi." All of yesterday and the first part of today was spent rebuilding the computer. Don't trust facebook. I now have a centos operating system, so I'll be learning to use Linux fast (hopefully).
  144   Sat Jun 27 22:09:36 2009 DmassLaserPSLI did some things in lab, some of which were bad.

I fired up the 35W, and everything was still working as per my last entry - we still could lock the PMC, and were generating small amounts of green.

  • I replaced the old home made mode matching lens slider mounts with real slider mounts
  • I then realized that I had no transverse adjustment with these sliders, which ended up being a royal PITA for alignment
  • I went back to the old mounts and tried to restore their position very faithfully, but had not documented their placement
  • I am now not aligned well into the PMC, and can't find the 00 - some combination of lens position adjustment and angular alignment should fix this.

 

I took a series of beam scans to find the waist size/position of the beam coming into the PMC and noticed a few things:

  • The beam scan IS UNHAPPY <edit with specific error> when it boots up the software
  • This *seems* to be making a number of features unavailable (i.e. the position dot, the contour plots, dual axis display...)
  • All of the things that I was unable to get to work involved needing both axes of the beamscan to work.
  • It seems like there is at least one ghost beam which is ~ on axis. I am unsure as to whether this is modal content of the PSL or a ghost from some optic.
  • The horizontal axis was FUGLY in modal content; the near vertical was pretty, and resembled a guassian. <Pics to be attached>

NEXT:

  • Troubleshoot: Look at the beamscan documentation regarding its startup errors/running the features that were "greyed out"
  • Enter data into mode matching program - get desired adjustments to lens positions
  • Verify orientation of lenses (both are plano convex, and I currently have the curved surfaces pointed towards each other)
  • Once I get a lock to the 00 mode, do fine tune of mode matching and angular alignment by hand
  • Move on to making my SHG more efficient, and maybe move up to tens of uW

 

  143   Fri Jun 26 17:20:45 2009 Michelle Stephens, Connor MooneyLaserFiberNPRO beam width divergence measurements

We used the BeamScan to characterize the divergence of the NPRO laser as a function of distance. Measurements were taken at (1/e^2)*Imax. We then fit the data to the function characteristic of gaussian divergence to find the location and size of the waist. The results are shown graphically in the attached file. Extrapolating, we found that the waist is 21.5 cm behind the laser aperture, and the (virtual, assuming that the laser beam encounters a diverging lens before the aperture) waist size is 124 micrometers.

 

Here is our data:

The distances, measured in inches from the base of the laser, are

[0 0.5 1 3 5 7 9 11 13 15 17 19 21 23 25]. The distance from the aperture of the laser to the base is 0.75 inches.

The beam sizes, measured in micrometers along one axis, are

[1334 1399 1542 1655 1968 2234 2537 2785 3063 3359 3638 4001 4253 4561 4708]

 

The directories for each of the Matlab files used to generate a graph of the data and carry out a least-squares regression are

/users/cmooney/BeamScans.m

/users/cmooney/myfun.m

Attachment 1: NPRObeamdiv.pdf
NPRObeamdiv.pdf
  142   Thu Jun 25 10:44:46 2009 MichelleMiscGeneralWeekly Summary 1

Over the course of the first week here I've done several things. The first thing we're going to try to do in the lab (once we have all our parts in) is lock a simple cavity using Pound-Drever-Hall. I've been reading some of the literature on that and familiarized myself with that technique. Through these readings, I've come across various terms and equipment used in optical setups that I was unfamiliar with, so I've become acquainted with those. I attended the laser safety training meeting and have started to find my way around the lab. Connor and I also began characterization of the Nd-YAG laser. While I will not be working with that one, I will have to do the same thing for the one I will be working with. We first made some measurements of its output power versus the drive current, and results from that can be found in an entry to the eLog on 6-23-09. Today we are going to use the BeamScan to figure out the divergence of the laser.

Various other things that I've done over the past week include learning how to work with Matlab and Linux (both of which I am relatively unfamiliar with), and attending talks about the LIGO project by Dr. Weinstein.

  141   Wed Jun 24 15:17:59 2009 DmassComputingCDSNokia Wireless Device

To use the small Nokia wireless device:

 

1) Connect to cdsrana

2) Check what the ip of the machine you want to use is (/sbin/ifconfig)   - 131.215.113.56 at the time of this entry

3) run an x-terminal on the nokia (under extras)

4) enter into the command line :   ssh controls@IP

5) It will prompt you for a password, enter.

you can now run simple medm screens of your design from anywhere with wireless access to the cdsrana network!

ELOG V3.1.3-